OpenStudy (anonymous):

how do I write this is sigma notation? infinite geometric series where a1 = 42 and the common ratio is three fourths.

OpenStudy (mosaic):

\[\Large 42\sum_{n = 1}^{\infty}(\frac 34)^{n-1}\]

OpenStudy (campbell_st):

well its almost right try \[\sum_{n = 1}^{\infty} 42 \times (\frac{3}{4})^{n - 1}\] n = 1 you get the 1st term 42

OpenStudy (anonymous):

one more thing it's divergent right? because its infinite

OpenStudy (campbell_st):

its convergent because the common ratio is between -1 < r < 1 when that occurs you have a convergent sequence